What is the electric field strength at a point 40 cm from a charge of 5x10⁹ C?

1
Expert's answer
2022-03-09T12:30:46-0500

Electric field

"E=\frac{kq}{r^2}"

Put value


"E=\frac{9\times10^9\times5\times10^9}{0.4^2}=2.82\times10^{20}N/C"